An idea] gas described by T_i = 300.0 K, P_i = 1.1
Answer #1

when heated at constant volume, work done= 0 ( for constant volume process work done =0)

Temperature during constant volume process can be calculated from

T2= P2*T1/P1= 9.8*300/1.1 =2673K

during isothermal expansion, the pressure changes to 1.1 bar from 9.8 bar(P3) and P2V2= P3V3

V3= P2V2/P3= 9.8*10/1.1=89.09L

work done during isothermal expansion = -RT ln(V2/V3)= -RTln (9.8/1.1)= -48622.19 joules

it is then restored to original state by constant pressure and work done = 1.1*(10-89.09) bar.=-8699.9 joules

total work = =48622.19-8699.9=-57322.09 joules

When the cycle were run in opposite direction ,work done remain at 57322.09 joules